xx TimerInfo

TimerInfo

Abstract

1 Introduction 2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11 clock 10 Scope: private01  STRING

11 Description: Which clock to use

11 Range 10              01  Default: all
11 .*  10 Anything

11

11 every 10 Scope: private01  INT

11 Description: How often to print information

11 Range 10              01  Default: (none)
11 0:*  10 Positive

11

11 timer 10 Scope: private01  STRING

11 Description: Which timer to use

11 Range 10              01  Default: all
11 .*  10 Anything

11

7 Interfaces

General Implements: timerinfo

8 Schedule This section lists all the variables which are assigned storage by thorn CactusExamples/TimerInfo. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxPOSTSTEP x timerinfo x print information from cactus timers x

11 x 11 Language: 11 c
11 x 11 Type:  11 function